Toil & Oil processes crude oil to jointly produce gasoline, diesel, and kerosene. One batch produces 3,415 gallons of gasoline, 2,732 gallons of diesel, and 1,366 gallons of kerosene at a joint cost of $12,000. After the split-off point, all products are processed further, but the estimated market price for each product at the split-off point is as follows:

Gasoline $2 per gallon
Diesel 1 per gallon
Kerosene 3 per gallon

Using the market value at split-off method, allocate the $12,000 joint cost of production to each product.

Joint Product Allocation
Gasoline $
Diesel
Kerosene
Totals $
Answer #1
Job Products Quality (per gallon) (i) Selling price ( per gallon) (ii) Sales Value (i) x(ii) Allocation rate Cost to be allocated
Gasoline 3,415 2 6,830 6,830/13,660= 50% 12,000 x 50% = 6,000
Diesel 2,732 1 2,732 2,732/13,660 = 20% 12,000 x 20% = 2,400
Kerosene 1,366 3 4,098 4,098/13,660 = 30% 12,000 x 30% = 3,600
Total $13,660 $12,000
Joint Product Allocation
Gasoline 6,000
Diesel 2,400
Kerosene 3,600
Totals $12,000

Kindly comment if you need further assistance. Thanks‼!